# HKCanto-Eval: A Benchmark for Evaluating Cantonese Language Understanding and Cultural Comprehension in LLMs

### v1

@@ Abstract
Tsz Chung Cheng1, Chung Shing Cheng2, Chaak Ming Lau3, Eugene Tin-Ho Lam4, Chun Yat Wong5, Hoi On Yu6, Cheuk Hei Chong7, 1 Kyushu University, 2 hon9kon9ize, 3 The Education University of Hong Kong, 4 Independent Researcher, 5 The University of Hong Kong, 6 Independent Researcher, 7 Votee Limited, Correspondence: jed.cheng@mag.ed.kyushu-u.ac.jp
@@ 1 Introduction
Recent advancements in large language models (LLMs) such as GPT-4, Gemini, and various open-weight models have demonstrated remarkable capabilities in natural language understanding across multiple languages (Xu et al., 2024). However, most models’ performances significantly decline when applied to languages other than English, yielding particularly poor outcomes for low-resource languages (LRLs). Many of these LRLs are under-represented lingua francas that play a crucial role in daily interactions that are poorly supported due to the lack of training data, which can be mitigated by leveraging nearby non-LRLs, and present low-hanging fruits for few-shot learning. Notable examples include regional languages in Indonesia (Aji et al., 2022; Winata et al., 2022). In the spirit of language sustainability and AI support for marginalised communities (Du et al., 2020), it is imperative to improve multilingual support for LRLs by creating benchmarks to guide the future development of multilingual LLMs. {CJK*}UTF8bsmi This paper investigates the status of LLM support for Cantonese (ISO 639-3: yue), a member of the Sinitic (“Chinese”) branch of the Sino-Tibetan language family, and a distinct variety unintelligible to users of Mandarin, the standard variety of Chinese used in Mainland China (Pǔtōnghuà) and Taiwan (Guóyǔ). Cantonese, spoken by over 85 million people according to Ethnologue (Eberhard et al., 2024), serves as the most common and de facto official language of Hong Kong and Macau, and is also widely used in parts of Guangdong, Guangxi, Malaysia, and Singapore. Additionally, it is used as a diasporic language in countries such as Canada (Sachdevl et al., 1987), the United States (Leung and Uchikoshi, 2012), Australia (Zhang et al., 2023), and the United Kingdom (Bauer, 2016; Tsapali and Wong, 2023). Despite its widespread use, Cantonese is still considered a low-resource language (Xiang et al., 2024) due to the lack of quality written resources. This scarcity results from a “diglossia" that requires Written Chinese (which resembles Mandarin) to be used in formal settings [^1] Even in Mandarin-like Written Chinese, there are persistent lexical differences with other regions due to vastly different governmental, legal and education systems. For instance, the word “taxi” is rendered as “[Uncaptioned image][Uncaptioned image][Uncaptioned image]” in mainland China, “[Uncaptioned image][Uncaptioned image][Uncaptioned image]” in Taiwan, and “[Uncaptioned image][Uncaptioned image]” in Hong Kong and Macau., and a longstanding, ideologically-driven stigmatisation of Cantonese as an informal/vulgar language (Lau, 2024), further confining written Cantonese to informal contexts like social media and texting.
@@
Cantonese is partially supported by certain LLMs, with models like GPT-4 and Gemini capable of comprehending and responding in Cantonese (Fu et al., 2024; Hong et al., 2024; Jiang et al., 2024). There are models dedicated to better supporting Chinese languages and dialects: The Hong Kong government is developing an internal tool based on locally developed LLMs for administrative use (Yiu, 2024); SenseTime released SenseChat (Cantonese) model trained on 6 billion tokens of Hong Kong-specific data (SenseTime, 2024). However, the current support level is mostly contributed to by small pockets of Cantonese presented in the sheer volume of Written Chinese training data, and the computational power used to train the models. The following section outlines how current benchmarking studies have yet to provide a comprehensive evaluation for Cantonese and Hong Kong-related tasks that taps into the in-depth representation of underlying aspects of the language, which we believe is the prerequisite for accurate comprehension in uncommon scenarios.
@@ 2 Related Benchmarks
Although most existing LLM benchmarks focus on English-language tasks, culturally-aware datasets integrating machine-translated questions, native datasets and exam questions have been developed in other languages, including Arabic (Koto et al., 2024), Spanish (Plaza et al., 2024), Indic languages (Verma et al., 2024), and Korean (Son et al., 2024). Similar benchmarks have been published for Chinese, such as CMMLU (Li et al., 2023) and C-Eval (Huang et al., 2024) that gathered questions from various academic and professional exams in mainland China, and TMLU (Chen et al., 2024) and TMMLU+ (Tam et al., 2024) that evaluate knowledge in Traditional Chinese in the context of Taiwan.
These benchmarks are not applicable to the Hong Kong context due to the aforementioned diglossia and regional lexical differences. Recently, Jiang et al. (2024) introduced a Cantonese evaluation benchmark that combines four datasets translated from other languages (ARC, GSM8K, CMMLU, and Truthful-QA) [^2] It also contains a translation evaluation component for English-Cantonese and Simplified-to-Traditional Chinese translations but its data sources and evaluation methods are not fully transparent., resulting in a dataset that is heavily biased towards American culture (16.9% entries in the Truthful-QA dataset reference the United States) or mainland Chinese exams (CMMLU) (see Appendix A).
@@ 3.1 Translated MMLU Dataset
The first dataset comprises 14,042 questions from the original MMLU dataset in English (Hendrycks et al., 2020) and their Cantonese translation333using the Google Gemini 1.5 Flash API, which offers a balance between top performance and cost as one would find in the later section. This allows us to compare how LLMs perform when handling knowledge in a wide range of subjects in Cantonese rather than in English (See Appendix B).
@@ 3.2 Academic and Professional Dataset
The Academic and Professional Dataset is a set of multiple-choice questions curated to measure LLMs’ reasoning and problem-solving abilities in domain-specific knowledge. The dataset contains multiple-choice questions from 3 sub-categories: (1) Academic: Questions sourced from Hong Kong Diploma of Secondary Education (HKDSE), a territory-wide high-school graduate-level exam; extracted from scanned PDFs and are believed to have never appeared online in a plain-text form; (2) Professional: Questions from seven professional qualification exams, extracted from text PDF files found on the corresponding official sites (in which the model answers were not on the same page as the questions, avoiding data contamination concerns), and an additional set of Taxi Licensing Exam Styled Route Planning questions on Hong Kong roads and geographical features; (3) Law: Questions about law in Hong Kong across 15 categories sourced from the Internet, and an additional subset of the Basic Law edited by the authors and included in this sub-category.
@@ 3.3 Hong Kong Cultural Questions Dataset
This dataset contains 277 manually crafted questions divided into five categories that capture cultural knowledge common to people who have lived or grown up in Hong Kong that are often not learned in schools. The categories are Food Culture, History and Landmarks, Language and Expressions, Life in Hong Kong and Local Area Knowledge. The questions were collected in a way to capture knowledge from all walks of life. 244 questions were developed by the authors and volunteers for the first four categories, and the last category comes from an online quiz. Questions were created so that they were non-trivial and at the same time not too obscured. Details can be found in Appendix D.
@@ 3.4.3 Grapheme-to-Phoneme (G2P) Conversion
This dataset addresses the task of converting a string of written text represented in Traditional Chinese characters into Jyutping, a widely adopted romanisation standard of Cantonese [^4] https://lshk.org/jyutping-scheme. This is similar to typical G2P tasks except that Jyutping is used instead of the International Phonetic Alphabet (IPA) as the output. G2P functionalities have been implemented by PyCantonese (Lee et al., 2022), a Cantonese NLP package, Hambaanglaang Converter [^5] https://test.hambaanglaang.hk and Visual Fonts [^6] https://visual-fonts.com. There are so far no reliable converters for the task. This part of the dataset contains 150 pairs of Character-Jyutping sentences from both Standard Written Chinese and Cantonese and in a range of formality levels, manually checked by professional linguists from the Linguistic Society of Hong Kong, the organisation that established and maintains the Jyutping system. The score calculation method is discussed in Appendix E.3.
@@ 3.7 Model Selection
Popular multilingual open-weight models including Aya 23 8B (Aryabumi et al., 2024), Gemma 2 2B, 9B and 27B (Gemma Team et al., 2024), Llama 3.1 8B, 70B and 405B (Dubey et al., 2024), and Mistral Nemo Instruct 2407 12B (Mistral, 2024) were included to assess their cross-lingual ability. The collection also included two open-weight multilingual models from Chinese companies, Yi 1.5 6B, 9B and 34B (Young et al., 2024) and Qwen2 7B and 72B (Yang et al., 2024). In addition, CLLM777anonymised for the review 6B and 34B are two of the few open-weight models trained specifically on Cantonese data. They were trained by fine-tuning Yi 1.5 6B and 34B models with around 400 million tokens of Hong Kong-related content. Open-weight instructions fine-tuned models smaller than 70B parameters were evaluated using Nvidia H100 GPUs. The 70B and 405B models were evaluated using the API of SiliconFlow [^8] https://siliconflow.cn.
@@ 4.2 Academic and Professional Questions
The results of this dataset showed expected problem-solving abilities across models in different subject areas, in particular, general weaknesses in handling secondary school-level mathematics and strong performance in legal questions. Proprietary models generally performed better than open-weight models. The sub-scores in the individual tasks show that most models struggled with academic questions that were never posted online. It is worth noting that some open-weight models (e.g. CLLM 34B and Qwen2 72B) outperformed most models, likely due to these models being trained on Hong Kong data that represents local knowledge. Written Chinese yielded better overall results, and this is contributed by the Law dataset which only came in Chinese. Discounting this set, Written Chinese caused a slight drop in performance. This indicates that multi-lingual open-weight LLMs showed cross-lingual capabilities, maintaining similar performance across both languages.
@@ 4.3 Hong Kong Cultural Questions
Proprietary models and Qwen 2 72B showed a good understanding of Hong Kong cultural knowledge, yet none of the models performed well across the subcategories. Looking into the sub-scores, models occasionally matched humans in most sub-tests (e.g. Food Culture and Life in HK ). However, when inspecting the results, good performance by percentage only reflects the size of existing Hong Kong knowledge represented in Wikipedia entries. For example, only two models (Yi 1.5 6B and Qwen2 72B) correctly answered the origin of Demae Itcho noodles sold in Hong Kong, while 94% of humans did. The results for Language & Expressions also show that most models did not have a nuanced understanding of Cantonese. Compared to human performance at 85.8%, SenseChat scored the highest point out of all models in 5-shot (79.6%), but its performance dropped significantly in zero-shot (61.4%). In zero-shot evaluation, CLLM 34B delivered the best performance at 77.3%. Furthermore, model size affects the performance of geospatial tasks, with open-source models in the 6-9B parameter range achieving only about 50% of larger models’ performance on Local Area Knowledge (e.g. Yi 1.5 34B 67.9%, 9B 35.7%). The overall results of this dataset suggest that Hong Kong cultural knowledge is under-represented in LLM training. See Appendix C for details.

### v2

@@ Abstract
Tsz Chung Cheng1, Chung Shing Cheng2, Chaak Ming Lau3, Eugene Tin-Ho Lam5, Chun Yat Wong4, Hoi On Yu5, Cheuk Hei Chong6,7 1 Kyushu University, 2 hon9kon9ize, 3 The Education University of Hong Kong, 4 The University of Hong Kong, 5 Independent Researcher, 6 Votee AI, 7 Beever AI Correspondence: Tsz Chung Cheng: jed.cheng@mag.ed.kyushu-u.ac.jp, Chung Shing Cheng: joseph.cheng@hon9kon9ize.com Chaak Ming Lau: lchaakming@eduhk.hk
@@ 1 Introduction
Recent advancements in large language models (LLMs) such as GPT-4, Gemini, and various open-weight models have demonstrated remarkable capabilities in natural language understanding across multiple languages (Xu et al., 2024). However, the performance of most models significantly declines when applied to languages other than English, yielding particularly poor outcomes for low-resource languages (LRLs). These languages are under-represented lingua francas that play a crucial role in certain communities, and it is imperative to improve multilingual support for LRLs by creating benchmarks to guide the future development of multilingual LLMs. Since they are poorly supported due to the lack of training data, if there is a close language with more resources, this problem can potentially be mitigated through few-shot learning. A notable example of this strategy is the use of Bahasa Indonesian to handle regional languages in Indonesia (Aji et al., 2022; Winata et al., 2022). This strategy aligns with the spirit of language sustainability and AI support for marginalised communities (Du et al., 2020), which is also applicable to Cantonese.
@@
This paper investigates the status of LLM support for Cantonese (ISO 639-3: yue), a member of the Sinitic (“Chinese”) branch of the Sino-Tibetan language family, and a distinct variety unintelligible to users of Mandarin, the standard variety of Chinese used in Mainland China (Pǔtōnghuà) and Taiwan (Guóyǔ). Cantonese, spoken by over 85 million people according to Ethnologue (Eberhard et al., 2024), serves as the most common and de facto official language of Hong Kong and Macau, and is also widely used in parts of Guangdong, Guangxi, Malaysia, and Singapore. Additionally, it is used as a diasporic language in countries such as Canada (Sachdevl et al., 1987), the United States (Leung and Uchikoshi, 2012), Australia (Zhang et al., 2023), and the United Kingdom (Bauer, 2016; Tsapali and Wong, 2023). Despite its widespread use, Cantonese is still considered a low-resource language (Xiang et al., 2024) due to the lack of quality written resources. This scarcity results from a “diglossia" that requires Written Chinese (which resembles Mandarin) to be used in formal settings [^1] Even in Mandarin-like Written Chinese, there are persistent lexical differences with other regions due to vastly different governmental, legal and education systems. For instance, the word “taxi” is rendered as “[Uncaptioned image][Uncaptioned image][Uncaptioned image]” in mainland China, “[Uncaptioned image][Uncaptioned image][Uncaptioned image]” in Taiwan, and “[Uncaptioned image][Uncaptioned image]” in Hong Kong and Macau., and a longstanding, ideologically-driven stigmatisation of Cantonese as an informal/vulgar language (Lau, 2024), further confines written Cantonese to informal contexts like social media and texting.
@@
Cantonese is partially supported by certain LLMs, with models like GPT-4 and Gemini capable of comprehending and responding in Cantonese (Fu et al., 2024; Hong et al., 2024; Jiang et al., 2024). There are models dedicated to better supporting Chinese languages and dialects: The Hong Kong government is developing an internal tool based on locally developed LLMs for administrative use (Yiu, 2024); SenseTime released SenseChat (Cantonese), a model trained on 6 billion tokens of Hong Kong-specific data (SenseTime, 2024). However, the current support level is mostly contributed to by small pockets of Cantonese presented in the sheer volume of Written Chinese training data. There have been comparisons between Chinese and Western models on how well languages spoken in China are handled (Wen-Yi et al., 2025), showing that Chinese models outperformed Western ones on Mandarin, but the same cannot be said for Cantonese or other languages in China. The following section outlines how current benchmarking studies have yet to provide a comprehensive evaluation for Cantonese and Hong Kong-related tasks that tap into the in-depth representation of underlying aspects of the language, which we believe is the prerequisite for accurate comprehension in uncommon scenarios.
@@ 2 Related Benchmarks
Although most existing LLM benchmarks focus on English-language tasks, culturally-aware datasets integrating machine-translated questions, native datasets, and exam questions have been developed in other languages, including Arabic (Koto et al., 2024), Basque (Etxaniz et al., 2024a, b), Spanish (Plaza et al., 2024), Indic languages (Verma et al., 2024), and Korean (Son et al., 2024). Similar benchmarks have been published for Chinese, such as CMMLU (Li et al., 2023) and C-Eval (Huang et al., 2024) that gathered questions from various academic and professional exams in mainland China, and TMLU (Chen et al., 2024) and TMMLU+ (Tam et al., 2024) that evaluate knowledge in Traditional Chinese in the context of Taiwan.
These benchmarks are not applicable to the Hong Kong context due to the aforementioned diglossia and regional lexical differences. Recently, Jiang et al. (2024) introduced a Cantonese evaluation benchmark that combines four datasets translated from other languages (ARC, GSM8K, CMMLU, and Truthful-QA) [^2] It also contains a translation evaluation component for English-Cantonese and Simplified-to-Traditional Chinese translations, but its data sources and evaluation methods are not fully transparent., resulting in a dataset that is heavily biased towards American culture (16.9% entries in the Truthful-QA dataset reference the United States) or mainland Chinese exams (CMMLU) (see Appendix A).
@@ 3.1 Translated MMLU Dataset
The first dataset comprises 14,042 questions from the original MMLU dataset in English (Hendrycks et al., 2020) and their Cantonese translation [^3] The translation was done by the Google Gemini 1.5 Flash API, which offers a balance between top performance and cost as one would find in the later section. To address concerns regarding the accuracy of LLM translation, we have selected 4 questions from each category for human checking. 202 out of 228 sentences were judged to be good by the raters.. This allows us to compare how LLMs perform when handling knowledge in a wide range of subjects in Cantonese rather than in English (See Appendix B).
@@ 3.2 Academic and Professional Dataset
The Academic and Professional Dataset is a set of multiple-choice questions curated to measure LLMs’ reasoning and problem-solving abilities in domain-specific knowledge. The dataset contains multiple-choice questions from 3 sub-categories: (1) Academic: Questions sourced from Hong Kong Diploma of Secondary Education (HKDSE), a territory-wide high-school graduate-level exam; extracted and manually corrected from scanned PDFs and are believed to have never appeared online in a plain-text form; (2) Professional: Questions from seven professional qualification exams, extracted from text PDF files found on the corresponding official sites (in which the model answers were not on the same page as the questions, avoiding data contamination concerns), and an additional set of Taxi Licensing Exam Styled Route Planning questions on Hong Kong roads and geographical features; (3) Law: Questions about law in Hong Kong across 15 categories sourced from the Internet, and an additional subset of the Basic Law edited by the authors.
@@ 3.3 Hong Kong Cultural Questions Dataset
This dataset contains 277 manually crafted questions divided into five categories that capture cultural knowledge common to people who have lived or grown up in Hong Kong, that are often not learned in schools. The categories are Food Culture, History and Landmarks, Language and Expressions, Life in Hong Kong and Local Area Knowledge. The questions were collected in a way to capture knowledge from all walks of life. 244 questions were developed by the authors and volunteers for the first four categories, and the last category comes from an online quiz. Questions were created so that they were non-trivial and at the same time not too obscure, and have been verified by all the authors. Details can be found in Appendix D.
@@ 3.4.3 Grapheme-to-Phoneme (G2P) Conversion
This dataset addresses the task of converting a string of written text represented in Traditional Chinese characters into Jyutping, a widely adopted romanisation standard of Cantonese [^4] https://lshk.org/jyutping-scheme. This is similar to typical G2P tasks except that Jyutping is used instead of the International Phonetic Alphabet (IPA) as the output. G2P functionalities have been implemented by PyCantonese (Lee et al., 2022), a Cantonese NLP package, Hambaanglaang Converter [^5] https://test.hambaanglaang.hk and Visual Fonts [^6] https://visual-fonts.com. As the task is non-deterministic, rule-based conversions are bound to be unreliable (although Visual Fonts have achieved very high accuracy now). There is also no reliable non-rule-based G2P system to our best knowledge. This part of the dataset contains 150 pairs of Character-Jyutping sentences from both Standard Written Chinese and Cantonese and in a range of formality levels, manually checked by professional linguists from the Linguistic Society of Hong Kong, the organisation that established and maintains the Jyutping system. The score calculation method is discussed in Appendix E.3.
@@ 3.6 Evaluation Method
The outputs of both translation and summarisation evaluation were evaluated and graded by paid undergraduate students and teaching assistants. The rubric can be found in Appendix F and G. As technology improves, future LLMs can perform the task to offer scalability. Nonetheless, the results from this human evaluation will be useful for verifying the validity and consistency of LLM-as-a-judge in the future.
@@ 3.7 Model Selection
Popular multilingual open-weight models including Aya 23 8B (Aryabumi et al., 2024), Gemma 2 2B, 9B and 27B (Gemma Team et al., 2024), Llama 3.1 8B, 70B and 405B (Dubey et al., 2024), and Mistral Nemo Instruct 2407 12B (Mistral, 2024) were included to assess their cross-lingual ability. The collection also included two open-weight multilingual models from Chinese companies, Yi 1.5 6B, 9B and 34B (Young et al., 2024) and Qwen2 7B and 72B (Yang et al., 2024). In addition, CantoneseLLM (CLLM) v0.5 6B and 34B [^7] https://huggingface.co/hon9kon9ize/CantoneseLLMChat-v0.5 are two of the few open-weight models trained specifically on Cantonese data. They were trained by fine-tuning Yi 1.5 6B and 34B models with around 400 million tokens of Hong Kong-related content. Open-weight instructions fine-tuned models smaller than 70B parameters were evaluated using Nvidia H100 GPUs. The 70B and 405B models were evaluated using the API of SiliconFlow [^8] https://siliconflow.cn.
@@ 4.2 Academic and Professional Questions
The results of this dataset showed expected problem-solving abilities across models in different subject areas, in particular, general weaknesses in handling secondary school-level mathematics and strong performance in legal questions. Proprietary models generally performed better than open-weight models. The sub-scores in the individual tasks show that most models struggled with academic questions that were never posted online. It is worth noting that some open-weight models (e.g. CLLM v0.5 34B and Qwen2 72B) outperformed most models, and we can conduct further investigation on what additional training data was used to achieve this performance. Written Chinese yielded better overall results, and this is attributed to the Law dataset, which only came in Chinese. Discounting this set, Written Chinese caused a slight drop in performance. This indicates that multi-lingual open-weight LLMs showed cross-lingual capabilities, maintaining similar performance across both languages.
@@ 4.3 Hong Kong Cultural Questions
Proprietary models and Qwen 2 72B showed a good understanding of Hong Kong cultural knowledge, yet none of the models performed well across the subcategories. Looking into the sub-scores, models occasionally matched humans in most sub-tests (e.g. Food Culture and Life in HK ). However, when inspecting the results, good performance by percentage only reflects the size of existing Hong Kong knowledge represented in Wikipedia entries. For example, only two models (Yi 1.5 6B and Qwen2 72B) correctly answered the origin of Demae Itcho noodles sold in Hong Kong, while 94% of humans did. The results for Language & Expressions also show that most models did not have a nuanced understanding of Cantonese. Compared to human performance at 85.8%, SenseChat scored the highest point out of all models in 5-shot (79.6%), but its performance dropped significantly in zero-shot (61.4%). In zero-shot evaluation, CLLM v0.5 34B delivered the best performance at 77.3%. Furthermore, model size affects the performance of geospatial tasks, with open-source models in the 6-9B parameter range achieving only about 50% of larger models’ performance on Local Area Knowledge (e.g. Yi 1.5 34B 67.9%, 9B 35.7%). The overall results of this dataset suggest that Hong Kong cultural knowledge is underrepresented in LLM training. See Appendix C for details.
Phonological Knowledge Orthographic Knowledge NLP
Model
Homo-
phone
Rhyme Misc.
Visual
Sim.
Canton.
Char.
Misc. Avg.
Claude 3.5 Sonnet 28.0% 64.0% 16.0% 50.0% 76.9% 59.3% 89.2%
Doubao Pro 16.0% 44.0% 16.0% 70.0% 80.8% 48.1% 87.0%
Ernie 4.0 28.0% 60.0% 18.0% 70.0% 80.8% 53.7% 82.7%
Gemini 1.5 Flash 12.0% 20.0% 24.0% 40.0% 73.1% 31.5% 83.2%
Gemini 1.5 Pro 16.0% 40.0% 24.0% 50.0% 88.5% 46.3% 87.9%
GPT4o 56.0% 96.0% 28.0% 50.0% 65.4% 63.0% 89.6%
GPT4o-mini 20.0% 60.0% 20.0% 30.0% 57.7% 40.7% 86.1%
SenseChat 16.0% 36.0% 22.0% 75.0% 76.9% 42.6% 78.8%
Aya 23 8B 12.0% 40.0% 14.0% 15.0% 19.2% 31.5% 70.1%
CLLM v0.5 6B 24.0% 8.0% 18.0% 20.0% 50.0% 27.8% 71.9%
CLLM v0.5 34B 28.0% 28.0% 14.0% 35.0% 76.9% 37.0% 73.3%
Yi 1.5 6B 28.0% 12.0% 12.0% 10.0% 50.0% 20.4% 56.6%
Yi 1.5 9B 36.0% 40.0% 24.0% 30.0% 57.7% 18.5% 72.2%
Yi 1.5 34B 16.0% 32.0% 26.0% 30.0% 61.5% 33.3% 82.9%
Gemma 2 2B 8.0% 24.0% 18.0% 25.0% 53.8% 22.2% 73.4%
Gemma 2 9B 20.0% 28.0% 24.0% 25.0% 50.0% 33.3% 85.0%
Gemma 2 27B 20.0% 12.0% 16.0% 25.0% 65.4% 24.1% 83.2%
Llama 3.1 8B 12.0% 16.0% 18.0% 25.0% 42.3% 38.9% 60.3%
Llama 3.1 70B 28.0% 40.0% 12.0% 30.0% 61.5% 35.2% 84.5%
Llama 3.1 405B 20.0% 44.0% 18.0% 35.0% 65.4% 50.0% 64.4%
Mistral Nemo 12B 12.0% 28.0% 10.0% 25.0% 23.1% 37.0% 68.8%
Qwen2 7B 8.0% 40.0% 12.0% 35.0% 46.2% 33.3% 66.8%
Qwen2 72B 12.0% 28.0% 16.0% 50.0% 76.9% 48.1% 83.5%
Random/Control 16.0% 28.0% 24.0% 30.0% 11.5% 27.8% 76.8%
Table 2: Model performance on Linguistic Knowledge Dataset multiple-choice questions and NLP tasks. The bottom row indicates the expected correctness from random selection for the Phonological and Orthographic Knowledge tasks. For NLP, the reported figure is the average evaluation of professionally prepared translations for translation tasks serving as a control.
Model CER Levenshtein
Claude 3.5 Sonnet 7.9% 0.018
Doubao Pro 20.9% 0.044
Ernie 4.0 34.4% 0.094
Gemini 1.5 Flash 34.7% 0.083
Gemini 1.5 Pro 15.3% 0.030
GPT4o 5.4% 0.009
GPT4o-mini 12.0% 0.023
SenseChat 54.4% 0.163
Aya 23 8B 96.6% 0.724
CLLM v0.5 6B 94.1% 0.859
CLLM v0.5 34B 23.4% 0.058
Yi 1.5 6B 99.0% 0.577
Yi 1.5 9B 97.2% 0.528
Yi 1.5 34B 79.6% 0.837
Gemma 2 2B 97.5% 0.524
Gemma 2 9B 73.0% 0.259
Gemma 2 27B 62.5% 0.201
Llama 3.1 8B 69.9% 0.270
Llama 3.1 70B 31.3% 0.086
Llama 3.1 405B 26.3% 0.074
Mistral Nemo 12B 59.8% 0.201
Qwen2 7B 97.3% 0.466
Qwen2 72B 74.0% 0.268
Rule Based 0.8% 0.001